OpenStudy (anonymous):

Alexei wants to hang a mirror in his boat but the mirror and frame must not have an area larger than 19.25 square feet. The mirror is 3 feet wide and 5 feet long. Which quadratic equation represents the area of the mirror and frame combined? (The border must have equal width on each side.)

OpenStudy (anonymous):

2x2 + 8x − 19.25 = 0 3x2 + 5x − 19.25 = 0 4x2 + 16x − 4.25 = 0 4x2 + 16x + 15 = 0

OpenStudy (john_es):

\[(2x+3)(2x+5)=4x^2+16x-4.25\]

OpenStudy (anonymous):

@John_ES so its b?

OpenStudy (john_es):

In the list you wrote is (c)